Between 1980 and 1990, a monthly rent for a one bedroom apartment increased by $20 per year. In 1987, the rent was $350 per month. Find an equation that gives the monthly rent in dollars, y, in terms of the year, t. Let t = 0, correspond to 1980.
Answer by ankor@dixie-net.com(22740)

Between 1980 and 1990, a monthly rent for a one bedroom apartment increased by $20 per year.
In 1987, the rent was $350 per month.
Find an equation that gives the monthly rent in dollars, y, in terms of the year, t. Let t = 0, correspond to 1980.
:
From the information given we can find the slope easily using "rise/run"
a $20 rise for each year, therefore m=20
:
let b = rent in 1980 (t=0)
y = 20t + b
In 1987 t=7, y=350, find b
20(7) + b = 350
140 + b = 350
b = 350 -140
b = 210
therefore the equation
y = 20t + 210